OpenStudy (scarlettfarra2000):

Which one of the following statements expresses a true proportion? A. 2 : 3 = 3 : 2 B. 3 : 5 = 12 : 20 C. 14 : 6 = 28 : 18 D. 42 : 7 = 6 : 2

jhonyy9 (jhonyy9):

try them step by step 2/3 = 3/2 - so i think not 3/5 = 12/20 how is this ?

jhonyy9 (jhonyy9):

try simplifie the 12/20 and see what will get

jhonyy9 (jhonyy9):

12/20 = (4*3)/(4*5) = ?

OpenStudy (scarlettfarra2000):

4*3=12 4*5=20

jhonyy9 (jhonyy9):

yes but after you simplifie the 4 from numerator with 4 from denominator what will get like result ?

OpenStudy (scarlettfarra2000):

\[\frac{ 3 }{ 5 }\]

jhonyy9 (jhonyy9):

so do you understand it now ?

OpenStudy (scarlettfarra2000):

Yes, thank you

jhonyy9 (jhonyy9):

was my pleasure good luck anytime than i m online bye